OpenStudy (anonymous):

Solve by COMPLETING THE SQUARE: \[3x^2-6x-5=0\]

OpenStudy (anonymous):

x^2-2x+1=5/3+1 (x-1)^2=8/5

OpenStudy (anonymous):

x-1=+/- sqrt(8/5) x=1+/- sqrt(8/5)

OpenStudy (anonymous):

yep. thanks for showing it